OpenStudy (anonymous):

Please help, for a test! Will give Medal! What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (sleepyhead314):

Sorry but we are not allowed to help on tests or quizzes... that would be considered cheating :/

OpenStudy (bradely):

(-5, 3) and (1, -3) m = (-3-3)/(1-(-5)) =-6/6=-1 y+3=-1(x-1) y+3=-x+1 x+y=-2 x/-2+y/-2 =1 x-intercept is =-2 Source: http://www.mathskey.com/question2answer/